# Muon Optimizes Under Spectral Norm Constraints

### v1

@@ 1 Introduction
Recent advances in optimization algorithms have shown promising potential to outperform traditional adaptive gradient methods in training large-scale neural networks [PQK24, LLCL24, LLH+24, JJB+24, YLW+24, VMZ+25, PL24, Li22, PXA+25, XZL+24]. A noteworthy example is the Lion optimizer [CLH+23], which was discovered through symbolic search and has demonstrated competitive empirical performance across diverse tasks despite its simple update rule. A theoretical foundation for Lion was established through the Lion-$\mathcal{K}$ framework [CLLL24], which generalizes Lion and unifies powerful optimization techniques such as mirror descent [KBB15, BT03], Nesterov momentum [SWLW23, Nes83], Hamiltonian descent [MPT+18], Frank–Wolfe algorithms [PXA+25, Jag13], and decoupled weight decay [LH19, LSY+25].
@@
We leverage the theoretical guarantees of Lion-$\mathcal{K}$ dynamics to establish convergence rates for Muon under both the deterministic (Theorem 2) and stochastic gradient (Theorem 3) implementations. We furthermore show that Muon with decoupled weight decay converges to a Karush–Kuhn–Tucker (KKT) point of the aforementioned implicit constrained optimization problem (Theorems 4 and 5).
@@ 2 Related work
Weight decay is a widely used regularization technique in deep learning, traditionally implemented as an $\ell_{2}$ penalty directly coupled with gradient-based parameter updates [KB15]. The concept of decoupled weight decay, introduced by AdamW [LH19], separates regularization from adaptive gradient computations. Empirical evidence suggests that this decoupling enhances training stability and improves generalization, making it a standard practice in modern adaptive optimizers [CLH+23, LSY+25]. Recently, [XL24] demonstrated that AdamW implicitly solves a constrained optimization problem given convergence. [CLLL24] proved that optimizers with bounded updates and decoupled weight decay inherently correspond to constrained optimization formulations, even without requiring convergence assumptions.
@@
Although several concurrent works have studied the convergence of Muon under various smoothness assumptions [LH25, ALP+25, Kov25, SHH+25], none of them consider Muon with decoupled weight decay, despite it being the variant that has demonstrated the most promising empirical results [LSY+25]. [SW25] analyzes Muon with decoupled weight decay through a Frank–Wolfe perspective, whereas our work is the first to use the Lion-$\mathcal{K}$ framework.
@@ 3 Main results
where $\mathbb{X}:=\mathbb{R}^{n\times m}$ is the space of real $n\times m$ matrices, $\mathcal{F}:\mathbb{X}\to\mathbb{R}$ is a differentiable loss function, and the expectation is taken over the data distribution $\mathcal{D}$ with independent and identically distributed samples $\xi\sim\mathcal{D}$. Given a realization of the function $\mathcal{F}(\mathbf{X},\xi)$, the stochastic gradient $\nabla\mathcal{F}(\mathbf{X},\xi)$ is defined as the gradient of $\mathcal{F}(\mathbf{X},\xi)$ with respect to the variable $\mathbf{X}$. We assume throughout the paper that $\nabla\mathcal{F}(\mathbf{X},\xi)$ has variance bounded by $\frac{\sigma^{2}}{n_{\mathrm{batch}}}$.
@@
where $\mathbf{X}$ represents the parameters, $\mathbf{M}$ and $\widetilde{\mathbf{M}}$ represent momentum, $\mathbf{G}$ is either the deterministic gradient $\nabla\mathcal{F}(\mathbf{X})$ or a stochastic gradient $\nabla\mathcal{F}(\mathbf{X},\xi)$, $\eta_{t}>0$ is the learning rate, $\beta_{1},\beta_{2}\in[0,1]$ are two momentum coefficients, $\lambda\geq 0$ is the weight decay coefficient, and $\operatorname{msgn}$ is known as the matrix sign function (see Definition 2).
@@
^$\widehat{\mathcal{F}}(\mathbf{X}):=\mathcal{F}(\mathbf{X})+\frac{1}{\lambda}\mathcal{K}^{*}(\lambda\mathbf{X}),$ (4)
@@
Despite the general Lion-$\mathcal{K}$ framework, Muon’s use of the nondifferentiable nuclear norm casts unique challenges in providing convergence guarantees. In this work, we provide an analysis tailored to Muon and rigorously establish that the iterates of (2) converge to a KKT point of (5).
@@
Proposition 1.
We then identify two Lyapunov functions that are used to verify convergence in terms of the conditions in Proposition 1. For the constraint condition $\left\lVert\lambda\mathbf{X}^{\star}\right\rVert_{\textup{op}}\leq 1$, we use the Lyapunov function
@@
which measures the distance from $\mathbf{X}$ to the constraint ball $\mathbb{B}:=\{\mathbf{X}\in\mathbb{X}\mid\left\lVert\lambda\mathbf{X}\right\rVert_{\textup{op}}\leq 1\}$. Following the update (2), we show in Proposition 2 that $\mathcal{V}_{\mathbb{B}}(\mathbf{X}_{t})$ decays exponentially fast when $\eta_{t}<\frac{1}{\lambda}$, i.e.
@@
Hence, $\mathbf{X}_{t}$ rapidly converges to $\mathbb{B}$ at a linear rate and never leaves it after entering, allowing us to use the second Lyapunov function
@@
where $c$ is an appropriately defined scalar.
Figure 1: Convergence behavior of the Muon optimizer. Although the primary objective value $\mathcal{F}(\mathbf{X})$ exhibits nonmonotonic fluctuations, the Lyapnuov functions $\mathcal{V}_{\mathbb{B}}$ and $\mathcal{V}_{\mathcal{K}}$ decrease monotonically within their respective domains — $\mathcal{V}_{\mathbb{B}}$ when the trajectory is outside $\mathbb{B}$, and $\mathcal{V}_{\mathcal{K}}$ once the trajectory enters $\mathbb{B}$.
Leveraging the two Lyapunov functions, we apply a generalization of LaSalle’s invariance principle for discrete-time stochastic processes and prove that under mild conditions, the dynamics of Muon converge to the set of KKT points of (5). Furthermore, we use $\mathcal{V}_{\mathcal{K}}(\mathbf{X},\mathbf{M})$ to show that
These convergence results are summarized by the following theorem.
Theorem 1 (Informal, see Theorems 4 and 5).
Assume $\eta_{t}<\frac{1}{\lambda}$, $\sum_{t=0}^{\infty}\eta_{t}=\infty$, and $\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty$. Under certain conditions, when $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2), we have that $\mathbf{X}_{t}$ converges to a KKT point of (5) a.s., regardless of initialization.
Precise statements and detailed proofs of these results can be found in Sections 6 and 7.
@@ 4 Preliminaries
[…]
@@
where $p\in[1,\infty]$. $\left\lVert\cdot\right\rVert_{p}$ is the entrywise $\ell_{p}$ norm, $\left\lVert\cdot\right\rVert_{\textup{tr}}$ is known as the trace norm or nuclear norm, $\left\lVert\cdot\right\rVert_{\textup{F}}$ is known as the Frobenius norm, and $\left\lVert\cdot\right\rVert_{\textup{op}}$ is known as the spectral norm. The dual norm of $\left\lVert\cdot\right\rVert_{p}$ is $\left\lVert\cdot\right\rVert_{q}$, where $\frac{1}{p}+\frac{1}{q}=1$, the dual norm of $\left\lVert\cdot\right\rVert_{\textup{tr}}$ is $\left\lVert\cdot\right\rVert_{\textup{op}}$, and $\left\lVert\cdot\right\rVert_{\textup{F}}$ is self-dual.
The following fact will be useful for our analysis.
@@
[…]
@@ 4.1 Assumptions
The stochastic samples $\xi\sim\mathcal{D}$ are independent and identically distributed (i.i.d.). Additionally, the stochastic gradient $\nabla\mathcal{F}(\mathbf{X},\xi)$ satisfies
$\mathbb{E}_{\xi\sim\mathcal{D}}[\nabla\mathcal{F}(\mathbf{X},\xi)]=\nabla\mathcal{F}(\mathbf{X})\quad\text{and}\quad\textup{Var}(\nabla\mathcal{F}(\mathbf{X},\xi))\leq\frac{\sigma^{2}}{n_{\mathrm{batch}}},$
where $\sigma^{2}$ is a finite constant and $n_{\mathrm{batch}}$ denotes the batch size.
Assumption 3 ensures that the variance of the stochastic gradient is uniformly bounded. We remark that Assumptions 2 and 3 are standard in the literature for the analysis of stochastic optimization algorithms, e.g. [BN24, DBBU22, LWC+24].
@@
$\mathcal{K}:\mathbb{X}\to\mathbb{R}$ is convex.
We remark that Assumption 4 implies that $\mathcal{K}$ is also closed and proper.
@@ 5 Background on Lion-$\mathcal{K}$
  • Decoupled weight decay $\lambda\mathbf{X}$, which reduces the parameter magnitude in addition to the update $\nabla\mathcal{K}(\widetilde{\mathbf{M}})$. This introduces a regularization effect (see Section 5.1) and is closely related to Frank–Wolfe style algorithms.
@@ 5.1 Effect of decoupled weight decay
^$\nabla\widehat{\mathcal{F}}(\mathbf{X}_{t})=\nabla\mathcal{F}(\mathbf{X}_{t})+\nabla\mathcal{K}^{*}(\lambda\mathbf{X}_{t})=\mathbf{0}.$
@@ 5.2 Lyapunov function for Lion-$\mathcal{K}$
It is not immediately obvious why the Lion-$\mathcal{K}$ ODE would serve to minimize ^$\widehat{\mathcal{F}}(\mathbf{X})$, as the ODE does not necessarily guarantee a monotonic decrease in ^$\widehat{\mathcal{F}}(\mathbf{X})$. However, we show in Appendix A that the Lion-$\mathcal{K}$ ODE minimizes the auxiliary function
@@
in the sense that it is monotonically decreasing along the ODE trajectories, i.e. $\frac{\textup{d}}{\textup{d}t}\mathcal{H}(\mathbf{X}_{t},\mathbf{M}_{t})\leq 0$ until a local minimum is achieved. In other words, $\mathcal{H}(\mathbf{X},\mathbf{M})$ admits a Lyapunov function of the Lion-$\mathcal{K}$ ODE (9). Here, $\mathcal{H}(\mathbf{X},\mathbf{M})$ is a joint function of position $\mathbf{X}$ and momentum $\mathbf{M}$. It can be interpreted as a Hamiltonian function in a physical metaphor, where ^$\widehat{\mathcal{F}}(\mathbf{X})$ is the potential energy, and the additional term represents a form of kinetic energy.
Moreover, minimizing $\mathcal{H}(\mathbf{X},\mathbf{M})$ is equivalent to minimizing the regularized objective ^$\widehat{\mathcal{F}}(\mathbf{X})$. This can be seen by the Fenchel–Young inequality, which ensures that $\mathcal{K}^{*}(\lambda\mathbf{X})+\mathcal{K}(\mathbf{M})-\left\langle\mathbf{M},\lambda\mathbf{X}\right\rangle\geq 0$, with equality when $\mathbf{M}\in\partial\mathcal{K}(\lambda\mathbf{X})$.
@@ 6.1 Spectral norm constraint
[…]

### v2

@@ 1 Introduction
Recent advances in optimization algorithms have shown promising potential to outperform traditional adaptive gradient methods in training large-scale neural networks [PQK24, LLCL24, LLH+24, JJB+24, YLW+25, VMZ+25, PL24, Li22, PXA+25, XZL+24]. A noteworthy example is the Lion optimizer [CLH+23], which was discovered through symbolic search and has demonstrated competitive empirical performance across diverse tasks despite its simple update rule. A theoretical foundation for Lion was established through the Lion-$\mathcal{K}$ framework [CLLL24], which generalizes Lion and unifies powerful optimization techniques such as mirror descent [KBB15, BT03], Nesterov momentum [SWLW23, Nes83], Hamiltonian descent [MPT+18], Frank–Wolfe algorithms [PXA+25, Jag13], and decoupled weight decay [LH19, LSY+25].
@@
We leverage the theoretical guarantees of Lion-$\mathcal{K}$ dynamics to establish convergence rates for Muon under both the deterministic (Theorem 3) and stochastic gradient (Theorem 4) implementations. We furthermore show that Muon with decoupled weight decay converges to the set of Karush–Kuhn–Tucker (KKT) points of the aforementioned implicit constrained optimization problem (Theorems 5 and 6).
@@ 2 Related work
Weight decay is a widely used regularization technique in deep learning, traditionally implemented as an $\ell_{2}$ penalty directly coupled with gradient-based parameter updates [LH19]. The concept of decoupled weight decay, introduced by AdamW [LH19], separates regularization from adaptive gradient computations. Empirical evidence suggests that this decoupling enhances training stability and improves generalization, making it a standard practice in modern adaptive optimizers [CLH+23, LSY+25]. Recently, [XL24] demonstrated that AdamW implicitly solves a constrained optimization problem given convergence. [CLLL24] proved that optimizers with bounded updates and decoupled weight decay inherently correspond to constrained optimization formulations, even without requiring convergence assumptions.
@@
Although several concurrent works have studied the convergence of Muon under various smoothness assumptions [LH25, ALP+25, Kov25, SHH+25], none of them consider Muon with decoupled weight decay, despite it being the variant that has demonstrated the most promising empirical results [LSY+25]. [SW25] analyzes Muon with decoupled weight decay through a Frank–Wolfe perspective, whereas our work is the first to use the Lion-$\mathcal{K}$ framework and to prove convergence with decreasing step sizes à la Robbins–Monro.
@@ 3 Main results
where $\mathbb{X}:=\mathbb{R}^{n\times m}$ is the space of real $n\times m$ matrices, $\mathcal{F}:\mathbb{X}\to\mathbb{R}$ is a differentiable loss function, and the expectation is taken over the data distribution $\mathcal{D}$ with independent and identically distributed samples $\xi\sim\mathcal{D}$. Given a realization of the function $\mathcal{F}(\mathbf{X},\xi)$, the stochastic gradient $\nabla\mathcal{F}(\mathbf{X},\xi)$ is defined as the gradient of $\mathcal{F}(\mathbf{X},\xi)$ with respect to the variable $\mathbf{X}$.
@@
where $\mathbf{X}_{t}$ represents the parameters, $\mathbf{M}_{t}$ and $\widetilde{\mathbf{M}}_{t}$ represent momentum, $\mathbf{G}_{t}$ is either the deterministic gradient $\nabla\mathcal{F}(\mathbf{X}_{t})$ or a stochastic gradient $\nabla\mathcal{F}(\mathbf{X}_{t},\xi_{t})$, $\eta_{t}>0$ is the learning rate, $\beta_{1},\beta_{2}\in[0,1)$ are two momentum coefficients, $\lambda\geq 0$ is the weight decay coefficient, and $\operatorname{msgn}$ is known as the matrix sign function (see Definition 2).
@@
$\widehat{\mathcal{F}}(\mathbf{X}):=\mathcal{F}(\mathbf{X})+\frac{1}{\lambda}\mathcal{K}^{*}(\lambda\mathbf{X}),$ (4)
@@
Despite the general Lion-$\mathcal{K}$ framework, Muon’s use of the nondifferentiable nuclear norm casts unique challenges in providing convergence guarantees. In this work, we provide an analysis tailored to Muon and rigorously establish that the iterates of (2) converge to the set of KKT points of (5).
@@
We can show that a point $\mathbf{X}$ is a KKT point if and only if the KKT score is zero, i.e., $\mathcal{S}(\mathbf{X})=0$, and the primal constraint $\left\lVert\mathbf{X}\right\rVert_{\textup{op}}\leq\frac{1}{\lambda}$ is satisfied. We then identify two Lyapunov functions that are used to verify convergence in terms of these conditions. For the constraint condition $\left\lVert\mathbf{X}\right\rVert_{\textup{op}}\leq\frac{1}{\lambda}$, we use the Lyapunov function
@@
which measures the distance from $\mathbf{X}$ to the constraint ball $\mathbb{B}:=\{\mathbf{X}\in\mathbb{X}\mid\left\lVert\lambda\mathbf{X}\right\rVert_{\textup{op}}\leq 1\}$. Following the update (2), we show that $\mathcal{V}_{\mathbb{B}}(\mathbf{X}_{t})$ decays exponentially fast when $\eta_{t}<\frac{1}{\lambda}$, i.e.
@@
Hence, $\mathcal{V}_{\mathbb{B}}$ converges to 0 at a linear rate, which implies that $\mathbf{X}_{t}$ rapidly converges to $\mathbb{B}$ and never leaves it after entering. Inside the ball, we use a second Lyapunov function
@@
where $c$ is an appropriately defined scalar. We show that Muon (approximately) monotonically decreases $\mathcal{V}_{\mathcal{K}}$ within the constraint set, which implies that the KKT score vanishes along the trajectory by a generalization of LaSalle’s invariance principle for discrete-time stochastic processes.
Figure 1: Convergence behavior of Muon. Although the primary objective value $\mathcal{F}(\mathbf{X})$ exhibits nonmonotonic fluctuations, the Lyapnuov functions $\mathcal{V}_{\mathbb{B}}$ and $\mathcal{V}_{\mathcal{K}}$ decrease monotonically within their respective domains — $\mathcal{V}_{\mathbb{B}}$ when the trajectory is outside $\mathbb{B}$, and $\mathcal{V}_{\mathcal{K}}$ once the trajectory enters $\mathbb{B}$.
Our main results are summarized by Figure 1 and the following theorems.
Theorem 1 (Informal, see Theorems 3 and 4).
When $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2) with $\eta_{t}=\eta=\Theta\left(\frac{1}{\sqrt{T}}\right)$ and $\textup{Var}(\mathbf{G}_{t})\leq\frac{\sigma^{2}}{n_{\mathrm{batch}}}$, $\min_{1\leq t\leq T}\mathbb{E}[\mathcal{S}(\mathbf{X}_{t})]=O\left(\frac{1}{\sqrt{T}}+\frac{\sigma}{\sqrt{n_{\mathrm{batch}}}}\right)$.
Theorem 2 (Informal, see Theorems 5 and 6).
Assume $\eta_{t}\leq\frac{1}{\lambda}$, $\sum_{t=0}^{\infty}\eta_{t}=\infty$, and $\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty$. Under certain conditions, when $\mathbf{X}_{t}$, $\mathbf{M}_{t}$, and $\widetilde{\mathbf{M}}_{t}$ are updated using (2), we have that $\mathbf{X}_{t}$ converges to the set of KKT points of (5) a.s., regardless of initialization.
Precise statements and detailed proofs can be found in Section 7.
@@ 4 Preliminaries
For a norm $\left\lVert\cdot\right\rVert$ on $\mathbb{X}$, and $r>0$, we let $\mathbb{B}_{\left\lVert\cdot\right\rVert}(r):=\{\mathbf{X}\in\mathbb{X}\mid\left\lVert\mathbf{X}\right\rVert\leq r\}$ denote the ball of radius $r$. When $\mathbb{D}\subseteq\mathbb{X}$ and the norm is clear from context, we let $d(\mathbf{X},\mathbb{D}):=\inf_{\mathbf{Y}\in\mathbb{D}}\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert$ denote the distance from $\mathbf{X}$ to $\mathbb{D}$. We let $\left\lVert\cdot\right\rVert_{*}$ denote the dual norm of $\left\lVert\cdot\right\rVert$, where
@@
where $\in[1,\infty]. p\left\lVert\cdot\right\rVert_{p}$ is the entrywise $\ell_{p}$ norm, $\left\lVert\cdot\right\rVert_{\textup{tr}}$ is known as the trace norm or nuclear norm, $\left\lVert\cdot\right\rVert_{\textup{F}}$ is known as the Frobenius norm, and $\left\lVert\cdot\right\rVert_{\textup{op}}$ is known as the spectral norm. The dual norm of $\left\lVert\cdot\right\rVert_{p}$ is $\left\lVert\cdot\right\rVert_{q}$, where $\frac{1}{p}+\frac{1}{q}=1$, the dual norm of $\left\lVert\cdot\right\rVert_{\textup{tr}}$ is $\left\lVert\cdot\right\rVert_{\textup{op}}$, and $\left\lVert\cdot\right\rVert_{\textup{F}}$ is self-dual.
@@
Let $\left\lVert\cdot\right\rVert$ be a norm on $\mathbb{X}$ with dual norm $\left\lVert\cdot\right\rVert_{*}$. If $\mathcal{K}(\mathbf{X})=\left\lVert\mathbf{X}\right\rVert$, then
@@ 4.1 Assumptions
The stochastic samples $\xi_{t}\sim\mathcal{D}$ are independent and identically distributed (i.i.d.). Additionally, the stochastic gradient $\nabla\mathcal{F}(\mathbf{X}_{t},\xi_{t})$ satisfies
$\mathbb{E}_{\xi_{t}\sim\mathcal{D}}[\nabla\mathcal{F}(\mathbf{X}_{t},\xi_{t})]=\nabla\mathcal{F}(\mathbf{X}_{t})\quad\text{and}\quad\textup{Var}(\nabla\mathcal{F}(\mathbf{X}_{t},\xi_{t}))\leq\frac{\sigma^{2}}{n_{\mathrm{batch}}},$
where $\sigma^{2}$ is a constant and $n_{\mathrm{batch}}$ denotes the batch size.
Assumption 4 (Iteration-wise bounded variance).
The stochastic gradient $\nabla\mathcal{F}(\mathbf{X}_{t},\xi_{t})$ satisfies
$\mathbb{E}_{\xi_{t}\sim\mathcal{D}}[\nabla\mathcal{F}(\mathbf{X}_{t},\xi_{t})]=\nabla\mathcal{F}(\mathbf{X}_{t})\quad\text{and}\quad\textup{Var}(\nabla\mathcal{F}(\mathbf{X}_{t},\xi_{t}))\leq\left[\frac{\sigma^{2}}{n_{\mathrm{batch}}}\right]_{t}.$
We remark that Assumptions 2 and 3 are standard in the literature for the analysis of stochastic optimization algorithms, e.g. [BWAA18, DBBU22, LWC+24, YLW+25].
@@
$\mathcal{K}:\mathbb{X}\to\mathbb{R}$ is convex. This implies that $\mathcal{K}$ is also closed and proper.
@@ 5 Background on Lion-$\mathcal{K}$
[…]
@@ 5.1 Effect of decoupled weight decay
$\nabla\widehat{\mathcal{F}}(\mathbf{X}_{t})=\nabla\mathcal{F}(\mathbf{X}_{t})+\nabla\mathcal{K}^{*}(\lambda\mathbf{X}_{t})=\mathbf{0}.$
@@ 5.2 Lyapunov function for Lion-$\mathcal{K}$
It is not immediately obvious why the Lion-$\mathcal{K}$ ODE would serve to minimize $\widehat{\mathcal{F}}(\mathbf{X})$, as the ODE does not necessarily guarantee a monotonic decrease in $\widehat{\mathcal{F}}(\mathbf{X})$. However, we show in Appendix A that the Lion-$\mathcal{K}$ ODE minimizes the auxiliary function
@@
in the sense that it is monotonically decreasing along the ODE trajectories, i.e. $\frac{\textup{d}}{\textup{d}t}\mathcal{H}(\mathbf{X}_{t},\mathbf{M}_{t})\leq 0$ until a local minimum is achieved. In other words, $\mathcal{H}(\mathbf{X},\mathbf{M})$ admits a Lyapunov function of the Lion-$\mathcal{K}$ ODE (9). Here, $\mathcal{H}(\mathbf{X},\mathbf{M})$ is a joint function of position $\mathbf{X}$ and momentum $\mathbf{M}$. It can be interpreted as a Hamiltonian function in a physical metaphor, where $\widehat{\mathcal{F}}(\mathbf{X})$ is the potential energy, and the additional term represents a form of kinetic energy.
Moreover, minimizing $\mathcal{H}(\mathbf{X},\mathbf{M})$ is equivalent to minimizing the regularized objective $\widehat{\mathcal{F}}(\mathbf{X})$. This can be seen by the Fenchel–Young inequality, which ensures that $\mathcal{K}^{*}(\lambda\mathbf{X})+\mathcal{K}(\mathbf{M})-\left\langle\mathbf{M},\lambda\mathbf{X}\right\rangle\geq 0$, with equality when $\mathbf{M}\in\partial\mathcal{K}(\lambda\mathbf{X})$.
@@ 6.1 Spectral norm constraint
For any update of the form $\mathbf{X}_{t+1}=\mathbf{X}_{t}+\eta_{t}(\mathbf{O}_{t}-\lambda\mathbf{X}_{t})$ with $\left\lVert\mathbf{O}_{t}\right\rVert\leq b$, $\eta_{t}\lambda\leq 1$, and $\lambda>0$, where $\left\lVert\cdot\right\rVert$ is any norm on $\mathbb{X}$ and $b$ is a constant, we have